Marvin's Taxi Service charges $2.00 for the first kilometer and $1.25 for each additional km. If the cab fare was $19.50, how far did the taxi go?
(A) 12 km (B) 13 km (C) 14 km (D) 15 km (E) 16 km

Answer by rfer(16322)

D) 15 km
allow for the first mile included in the $2.00
2.00+1.25x-1.25=19.50
combine like
.75+1.25x=19.50
subtract .75 from both sides
.75-.75+1.25x=19.50-.75
1.25x=18.75
x=15